OpenStudy (anonymous):

Factor fully: 3-48x^2

OpenStudy (nenadmatematika):

=3(1-16x^2)=3(1-4x)(1+4x)

OpenStudy (anonymous):

Difference of the squares

OpenStudy (anonymous):

\[12m^{2}-36m+27\]

OpenStudy (anonymous):

plz help

OpenStudy (amistre64):

what have you tried?

OpenStudy (amistre64):

do 12 36 and 27 have any factors in common?

OpenStudy (anonymous):

Nope (JK)

OpenStudy (anonymous):

yeah so far i have \[3(4m^{2}-12m+9)\]

OpenStudy (anonymous):

3-48x^2 = 3 (1 - 16x^2) = 3 [ 1 - (4x)^2 ] = 3 [ (1 - 4x) ( 1+ 4x) ]

OpenStudy (amistre64):

good; now from there I multiply first and last (4*9) to get 36 the + at the end tells me the factors have the same sign; and the - in the middle tells me they are both negative

OpenStudy (amistre64):

3(x-?/4)(x-??/4) what factors of 36 add up to get 12?

OpenStudy (anonymous):

nenad you can always enclose your answers withing \.[ \.] (without dots) to get a beautiful latex effect :) Example: 3(1-16x^2)=3(1-4x)(1+4x) compared to : \[ 3(1-16x^2)=3(1-4x)(1+4x) \]

OpenStudy (anonymous):

@FoolsForMath someone already answer it. Lol

OpenStudy (anonymous):

6 and 6

OpenStudy (anonymous):

x = 3/4 , tinydancer

OpenStudy (anonymous):

@Hershey_Kisses: I was addressing a different issue.

OpenStudy (amistre64):

right :) 1 36 2 18 4 9 3 12 6 6 .... 6+6 = 12 :) lets use those 3 (x-6/4) (x-6/4) reduce the fractions 3 (x-3/2) (x-3/2) and stick the denom in front 3 (2x-3) (2x-3) and the rest is just making it look better

OpenStudy (amistre64):

if this method doesnt work; that means we eighet got complex issues, or fractions to deal with

OpenStudy (amistre64):

eighet is the universal typo for either lol

OpenStudy (anonymous):

awesome thanks :)